# Unveiling Downstream Performance Scaling of LLMs: A Clustering-Based Perspective

### v2

@@ 1 Introduction
Despite extensive efforts, a reliable scaling law for downstream tasks remains elusive. One line of work attempts to extrapolate the performance of a large model by modeling the performance-loss relationship [6, 13, 8, 38, 28], but this often fails to capture the emergent behaviors of LLMs and the mismatch between the in-domain loss and downstream metrics [43]. Another line of research focuses on direct extrapolation of the performance-compute relationship [1, 20], yet the uneven difficulty distribution between different evaluation samples undermines its accuracy.
A critical oversight in current approaches is the assumption of uniform scaling behavior across diverse evaluation samples. We observe that different evaluation samples actually follow distinct performance scaling patterns, and thus applying a single extrapolation formula to the entire evaluation set is suboptimal. We give a detailed analysis in Sec. 3.
@@
We validate our COD approach on eight popular evaluation sets, including MATH [16], BBH [33], and MMLU pro [36] datasets. COD achieves an average prediction error of 1.36% on an LLM with 70B parameters. Our results demonstrate that this difficulty-aware framework substantially outperforms existing methods, establishing a promising paradigm for accurate downstream performance scaling of LLMs.
@@ 2.2 Downstream Task Performance Scaling
Predicting downstream task performance remains a critical challenge due to emergent abilities in LLMs that manifest only after exceeding task-specific thresholds [37, 29]. Recent works, such as using loss as a proxy [6] or increasing metric resolution [20], have demonstrated potential but encounter challenges in aligning surrogate metrics with original task objectives. Here, we briefly review the two main types of methods for predicting downstream performance:
@@
End-to-end performance-compute prediction. These methods [20, 28, 1] directly model the relationship between performance and compute (or the number of model parameters). Additionally, Achiam et al. [1] estimate and fit this relationship using a subset of the evaluation set, while still failing to predict the full set. Hu et al. [20] address the challenge of non-emergent capabilities in smaller models by employing multiple non-greedy decoding evaluations, thereby enabling accurate extrapolation of performance predictions for models with up to 2.4B parameters.
@@ 3 Pilot Study
Fig. 1(left) illustrates the performance–loss relationships for LLMs of different sizes on the CEval benchmark [30]. At the same training loss level, smaller models can outperform larger ones in terms of test accuracy. Because smaller models initially exhibit weaker in-domain fitting capacity, they typically require more training steps to reach the same loss value, which can lead to better in-domain generalization once they do. Fig. 1(middle) compares the performance of LLMs trained under different learning rate schedules on the GSM8k dataset [7]. At the same loss level, the performance under the cosine schedule is always worse than that under the constant schedule, indicating that a lower learning rate may prioritize memorization over generalization, thereby diminishing downstream performance.
Diverse scaling patterns within the evaluation set. Different task samples exhibit unique computational thresholds, learning slopes, and upper bounds, making it challenging to find a single fitting function (or function group) that generalizes well across diverse task samples. Fig. 1(right) illustrates the performance-compute relationships on three random clusters of the BBH benchmark [33], with each cluster containing samples with similar difficulty. Even within a single evaluation set, these scaling curves can vary significantly, indicating that a one-size-fits-all performance-compute curve is insufficient for capturing the full spectrum of a downstream evaluation set.
@@ 4 Method
In this section, we introduce the COD method in four parts, as illustrated in Fig. 2: 1) We show the advantages of COD and present an improved mean-shift clustering algorithm (Sec. 4.1); 2) We derive a performance scaling law corresponding to task difficulty variance, which supports extrapolating the performance-compute relationship for task clusters with similar difficulty features (Sec. 4.2). We fit cluster-wise performance-compute curves on small models and filter extrapolatable clusters; 3) We extrapolate the performance on extrapolatable clusters and predict the accuracy of the target large model on the predictable subset(Sec. 4.3); 4) We show how to map accuracy on the predictable subset to full evaluations (Sec. 4.4).
@@ 4.1 Clustering on Difficulty
Tasks within evaluation sets, despite common themes, show substantial difficulty differences, leading to diverse scaling patterns that hinder universal fitting functions. We propose clustering tasks by similar performance scaling behaviors to minimize intra-cluster heterogeneity while ensuring robust evaluation with sufficient samples per cluster. To quantify task difficulty, we leverage the passrate(mean accuracy across multiple trials) from small models of progressively increasing scales. This process generates a feature vector for each task, which ideally exhibits monotonic growth, thereby reflecting the incremental enhancement of model capabilities.
Improved clustering methods. We hope to adopt clustering algorithms with the following features: 1. Minimizing intra-class variance to ensure similar extrapolation properties within each cluster, 2. Automatic determination of cluster numbers, as the optimal number varies across evaluation sets and is difficult to pre-specify.
While K-Means [27] requires pre-specifying cluster numbers and DBSCAN [11] may yield large intra-cluster variance, MeanShift [12] allows constraining this variance via a radius parameter, better suiting our needs.
@@
To further reduce intro-class variance, we propose an improved MeanShift algorithm to constrain the cluster diameter. At the same time, we maintain a minimum number of tasks in each cluster to reduce metric fluctuations. We provide the t-SNE visualization of evaluation tasks on BBH [33]. Each point represents an evaluation sample, and its color denotes the cluster type. As shown in Fig. 3, our improved MeanShift effectively splits dense areas. We explain the details of clustering algorithms in Sec. A.1, and smoothing techniques in Sec. A.2.
@@ 4.2 Fitting
After clustering, we compute metrics for small models within each cluster. These small models are trained with the same ratio of training tokens to compute per token. We then apply a novel performance scaling law, based on theoretical analysis, to fit accuracy-to-compute curves for each cluster, focusing on clustered samples after excluding outliers. We derive the following fitting formula for the downstream task scaling law based on the training loss scaling law.
@@
Given a language model trained with computational budget $C$, and a set of downstream tasks $P$, under the following assumptions:
  1. 1.
    Given a task $(q,\mathrm{ans})\in P$, the relationship between the answer loss and the compute follows a power law, $\frac{1}{n}\sum_{(q,\mathrm{ans})}\mathrm{loss}_{\mathrm{ans}}(C)\sim aC^{-b}+c$;
  2. 2.
    For tasks with a finite answer set, the model gives a random guess choice if it cannot truly solve it, i.e. $\mathbb{E}[\mathrm{Acc|incorrect}]=g$;
  3. 3.
    Task passrate equals the product of the predicted probability of each token $p_{\mathrm{ans}}=\prod_{t\in\mathrm{ans}}p(t)$, which means that each task has a unique answer and model output answer only without thinking progress.
The expected accuracy on tasks $P$ can be modeled as:
$\mathbb{E}_{p}[\mathrm{Acc}(C)]=g+(1-g)\left(e^{-aC^{-b}-c}+\frac{\sigma^{2}}{\mu}\right)+o(\mu),$ (1)
where $g$ represents the random guess performance floor, $1-g$ represents the maximum achievable performance improvement, $a,b,c$ are positive constants, and $\mu=\frac{1}{\#P}\sum_{(q,\mathrm{ans})\in P}\mathrm{loss}_{\mathrm{ans}}$; $\sigma^{2}=\frac{1}{\#P}\sum(\mathrm{loss}_{\mathrm{ans}_{t}}-\mu)^{2}$.
@@
Assumption 3 (unique answers, no pre-answer reasoning) implies a negative logarithmic relationship between $\mathrm{loss}_{\mathrm{ans}}$ and individual task passrate, as correct task completion probability becomes the product of answer token probabilities. A key distinction, often overlooked, is how averaging occurs: the passrate metric averages $\exp(-\mathrm{loss}_{\mathrm{ans}})$ (an arithmetic mean), whereas applying the loss scaling law effectively averages loss before exponentiation (yielding a geometric mean). We show that the difference between these means can be approximated by $\frac{\sigma^{2}}{\mu}$ (where $\sigma^{2}$ and $\mu$ are the variance and mean of task passrates, respectively). Consequently, a performance scaling law derived from the loss scaling law is most valid for evaluation sets with limited difficulty variance. Our COD method achieves this by clustering tasks, thereby constraining intra-cluster difficulty variance and enabling better alignment with the performance scaling law. Finally, for tasks with a finite answer set, a random guessing score $g$ accounts for expected performance on unmastered tasks. Detailed proofs are provided in Appendix B. ∎
theorem 1 demonstrates that a metric of an evaluation set with similar difficulty features can be effectively modeled using the following formula:
@@
where $a$ and $b$ jointly influence how accuracy varies with $C$, $c$ controls the upper bound of the fitting curve, and $g$ represents the expected random guess metric for a task cluster. Parameters $a$, $b$, $c$, and $g$ are to be fitted. Note that these assumptions may not perfectly hold in practice, we provide additional discussions on assumption 3 in Appendix G.
@@ 4.3 Extrapolation
To ensure reliable extrapolation, we identify clusters exhibiting robust scaling patterns, as some clusters may show saturated or non-emergent performance on smaller models, making them unsuitable for prediction.
@@
We filter out non-extrapolatable clusters using two rules based on the parameters from Eq. 2:
@@
In practice, for extrapolatable clusters, we set $a>1$, $b>0.1$, and $10\leq c<$ 1.
@@ 4.4 Mapping from Predictable Subset to Target Evaluation Set
We map predictions from the predictable subset $T^{\prime}$ to the complete evaluation set $T$ using a function $f$. This approach is based on the observation that extrapolatable and non-extrapolatable samples, while differing in difficulty, often share question types, suggesting a preserved relative metric ordering. The mapping function $f:T^{\prime}\rightarrow T$ is continuous, smooth over $[0,1]$, monotonically increasing, and constrained to pass through $(0,0)$ and $(1,1)$. Empirical validation indicates a quartic function optimally captures this relationship:
$f(x)=\alpha_{1}x^{4}+\alpha_{2}x^{3}+\alpha_{3}x^{2}+(1-\alpha_{1}-\alpha_{2}-\alpha_{3})x$ (3)
To ensure reliability, we calibrate $f$ using evaluation results from existing models as anchors. This subset-to-full mapping generally demonstrates robustness across diverse model architectures and training data, often permitting the use of external models (e.g., Qwen2-72B [40]) as anchors for many tasks (see Sec. 5.3.3 for experiments). For data-sensitive tasks, however, models with similar training distributions provide more reliable anchors, prioritizing data consistency for mapping accuracy. The final metric prediction for a target LLM with estimated training computation $C_{0}$ is then $p=f\circ y(C_{0})$, combining the cluster-wise extrapolation $y(C_{0})$ from Eq. 2 with the mapping $f$.
@@ 5.1 Experimental Setups
In our experimental setup, we train several smaller versions of the target model architecture for prediction. These models vary in size, but share similar training procedures, with the training data scaled proportionally to their sizes. We show the training configurations of the model in Appendix C.
For evaluation, we adopt the following widely used benchmarks, including GSM8K [7], MATH [16], BBH [33], TriviaQA [21], MBPP [3], AGIEval [44], DROP [9], MMLU-pro [36]. All models are evaluated in a few-shot in-context learning manner, and we aligned our evaluation setups with LLaMa3 [10]. We evaluate our proposed COD performance scaling for LLMs against existing approaches on multiple public benchmarks. Using a series of smaller models with identical data distribution and architecture but varied configurations, we estimate the downstream task performance of a pretrained LLM with 70B parameters.

### v3

@@ 1 Introduction
Despite extensive efforts, a reliable scaling law for downstream tasks remains elusive. One line of work attempts to extrapolate the performance of a large model by modeling the performance-loss relationship [8, 16, 11, 43, 31], but this often fails to capture the emergent behaviors of LLMs and the mismatch between the in-domain loss and downstream metrics [49]. Another line of research focuses on direct extrapolation of the performance-compute relationship [1, 23], yet a single family of curves usually fails to capture the performance on evaluation benchmarks with complex difficulty distributions across samples.
A critical oversight in current approaches is the assumption of uniform scaling behavior across diverse evaluation samples. We observe that different evaluation samples actually follow distinct performance scaling patterns, and thus applying a single extrapolation formula to the entire evaluation set is suboptimal. We give a detailed analysis in Sec.˜3.
@@
We validate our COD approach on eight popular evaluation sets, including MATH [19], BBH [38], and MMLU pro [41] datasets. COD achieves an average prediction error of 1.36%ha on an LLM with 70B parameters. Our results demonstrate that this difficulty-aware framework substantially outperforms existing methods, establishing a promising paradigm for accurate downstream performance scaling of LLMs.
@@ 2.2 Downstream Task Performance Scaling
Predicting downstream task performance remains a critical challenge due to emergent abilities in LLMs that some capabilities manifest only after exceeding task-specific thresholds [42, 32]. Recent works, such as using loss as a proxy [8] or increasing metric resolution [23], have demonstrated potential but encounter challenges in aligning surrogate metrics with original task objectives. Here, we briefly review the two main types of methods for predicting downstream performance:
@@
End-to-end performance-compute prediction. These methods [23, 31, 1, 7] directly model the relationship between performance and the compute budget (or the number of model parameters). Specifically, Achiam et al. [1] estimate and fit this relationship using a subset of the evaluation set, while still failing to predict the full set. Hu et al. [23] address the challenge of non-emergent capabilities in smaller models by employing multiple non-greedy decoding evaluations, thereby enabling accurate extrapolation of performance predictions for models with up to 2.4B parameters. Caballero et al. [7] propose a smooth broken power-law that models LLM scaling by decomposing it into multi-segment power laws. However, when predicting metrics for large-scale models (e.g., 70B parameters), performance trends often exhibit unexpected inflection points due to emergent capabilities or saturation effects, making piecewise functions inadequate for capturing these novel scaling regimes.
@@ 3 Pilot Study
figure˜1(left) illustrates the performance–loss relationships for LLMs of different sizes on the CEval benchmark [33]. At the same training loss level, smaller models can outperform larger ones in terms of test accuracy. Because smaller models initially exhibit weaker in-domain fitting capacity, they typically require more training steps to reach the same loss value, which can lead to better in-domain generalization once they do. figure˜1(middle) compares the performance of LLMs trained under different learning rate schedules on the GSM8k dataset [9]. At the same loss level, the performance under the cosine schedule is always worse than that under the constant schedule, indicating that a lower learning rate may prioritize memorization over generalization, thereby diminishing downstream performance.
Diverse scaling patterns within the evaluation set. Different task samples exhibit unique computational thresholds, learning slopes, and upper bounds, making it challenging to find a single fitting function (or function group) that generalizes well across diverse task samples. figure˜1(right) illustrates the performance-compute relationships on three random clusters of the BBH benchmark [38], with each cluster containing samples with similar difficulty. Even within a single evaluation set, these scaling curves can vary significantly, indicating that a one-size-fits-all performance-compute curve is insufficient for capturing the full spectrum of a downstream evaluation set.
@@ 4 Method
In this section, we first introduce the problem formulation. Then we present the COD method in four parts, as illustrated in figure˜2: 1) We show the construction of sample-level difficulty scaling feature and present an improved mean-shift clustering algorithm (Sec.˜4.1); 2) We derive a performance scaling law corresponding to task difficulty variance, which supports extrapolating the performance-compute relationship for task clusters with similar difficulty features (Sec.˜4.2). We fit cluster-wise performance-compute curves on small models and filter extrapolatable clusters; 3) We extrapolate the performance on extrapolatable clusters and predict the accuracy of the target large model on the predictable subset (Sec.˜4.3); 4) We show how to map accuracy on the predictable subset to full evaluations (Sec.˜4.4).
Problem Formulation. Consider a language model $M_{C}$ trained with a compute budget of $C$ measured in FLOPs. Let $\mathcal{P}$ be a set of downstream tasks that we aim to evaluate the model on. Each sample $T\in\mathcal{P}$ is defined by a question-answer pair $(q,a_{\text{true}})$. Given a question $q$, the model $M_{C}$ outputs a probability distribution $p(a|q;M_{C})$ over the space of all possible answers.
Our goal is to predict the downstream task performance of a large language model $M_{C_{\text{target}}}$ using only evaluation results from smaller models $\{M_{C_{1}},M_{C_{2}},\ldots,M_{C_{n}}\}$ where $C_{i}\ll C_{\text{target}}$ for all $i$. Formally, we aim to found the prediction method $\phi$ to minimize the absolute prediction error over a group of tasks sets $\{\mathcal{P}_{j}\}_{m}$:
$\arg\min_{\phi}\frac{1}{m}\sum_{i=1}^{m}\frac{1}{|\mathcal{P}_{j}|}\sum_{T\in\mathcal{P}_{j}}|\widehat{\text{Acc}}(C_{\text{target}},T)-\text{Acc}(C_{\text{target}})|,$
$\widehat{\text{Acc}}(C_{\text{target}},T):=\phi\left(\{\text{Acc}(C_{i},T)\}_{i=1}^{n},\{C_{i}\}_{i=1}^{n},C_{\text{target}}\right),$
where $\text{Acc}(C,T)$ denotes the accuracy of model $M_{C}$ on task $T$, and $\widehat{\text{Acc}}(C_{\text{target}},T)$ is the predicted accuracy for the target model.
@@ 4.1 Clustering on Difficulty
Tasks within evaluation sets, despite common themes, show substantial difficulty differences, leading to diverse scaling patterns that hinder universal fitting functions. We propose clustering tasks by similar performance scaling behaviors to minimize intra-cluster heterogeneity while ensuring robust evaluation with sufficient samples per cluster.
Specifically, we train a group of language models with increasing parameter counts. These models are trained with the same ratio of training tokens to compute per token. For each task, we generate 100 samples using top_p=0.7 and temperature=1.0 for each model, and compute the pass rate by averaging the results. This pass rate serves as an estimate of the model’s expected accuracy on the task. The resulting values are organized into a difficulty vector, ordered by increasing model size. For most tasks, this difficulty vector exhibits a monotonic increase, reflecting the gradual improvement of model capability with scale.
@@
After obtaining the difficulty feature vector for each task, we use the improved clustering algorithm that incorporates the following features: (1) Minimizing intra-class variance to ensure similar extrapolation properties within each cluster; (2) Automatic determination of cluster numbers, as the optimal number varies across evaluation sets and is difficult to pre-specify.
To further reduce intra-class variance, we propose an improved MeanShift algorithm to constrain the cluster diameter. At the same time, we maintain a minimum number of tasks in each cluster to reduce metric fluctuations. We provide the t-SNE visualization of clustering results evaluation tasks on BBH [38] to compare the proposed method and classic clustering algorithms including DBSCAN [14] and MeanShift [15]. Each point represents an evaluation sample, and its color denotes the cluster type. As shown in figure˜3, our improved MeanShift effectively splits dense areas whereas DBSCAN and the original MeanShift produce connected clusters with large within-cluster distances.
We provide numerical comparison of clustering algorithms in Sec.˜5.3.1 and explain implementation details of imporved MeanShift in Appendix A.1, smoothing techniques in Appendix A.2.
@@ 4.2 Fitting
After clustering, we compute metrics for small models within each cluster. We then apply a novel performance scaling law, based on theoretical analysis, to fit accuracy-to-compute curves for each cluster, focusing on clustered samples after excluding outliers. We derive the following fitting formula for the downstream task scaling law based on the training loss scaling law.
@@
Consider a language model $M_{C}$ trained with compute budget $C$ and a set of downstream tasks $\mathcal{P}$. Under the following assumptions: Assumption 1 (Power-law scaling of answer loss): the expected answer loss follows:
$L_{\mathcal{P}}(C):=\mathbb{E}_{(q,a_{\text{true}})\sim\mathcal{P}}[L(q,a_{\text{true}};C)]=\alpha C^{-\beta}+\gamma$ (1)
where $\alpha,\beta,\gamma>0$ are task-specific constants, with $\gamma$ representing the irreducible loss.
Assumption 2 (Unique deterministic answers): Each question has a unique deterministic answer. The model receives score $1$ if and only if $M_{C}$ outputs $a_{\text{true}}$, and 0 otherwise.
Assumption 3 (Accuracy decomposition): The expected accuracy decomposes as:
$\mathbb{E}_{T\sim\mathcal{P}}[\mathrm{Acc}(C)]=g+(1-g)\cdot\mathbb{E}_{(q,a_{\text{true}})\sim\mathcal{P}}[p(a_{\text{true}}|q,M_{C})],$ (2)
where $g\in[0,1]$ is the random guessing baseline.
Then, the expected accuracy on task set $\mathcal{P}$ can be modeled as:
$\mathbb{E}_{\mathcal{P}}[\mathrm{Acc}(C)]=g+(1-g)\left(\exp{(-\alpha C^{-\beta}-\gamma)}+\frac{\sigma_{L}^{2}(C)}{2\mu_{L}(C)}\right)+o\left(\sigma_{L}^{2}(C)\right)$ (3)
where $\mu_{L}(C)=\mathbb{E}_{(q,a_{\text{true}})\sim\mathcal{P}}[L(q,a_{\text{true}};C)]$ is the mean loss and $\sigma_{L}^{2}(C)=\mathrm{Var}_{(q,a_{\text{true}})\sim\mathcal{P}}[L(q,a_{\text{true}};C)]$ is the loss variance across the task set.
@@
By definition of the language model loss, $p(a_{\text{true}}|q,M_{C})=\exp(-L(q,a_{\text{true}};C))$. Under Assumption 1, if the answer loss follows a power law $L\sim\alpha C^{-\beta}+\gamma$, then the task passrate should approximately scale as $\exp(-\alpha C^{-\beta}-\gamma)$.
The key subtlety lies in the averaging: accuracy computes $\mathbb{E}[\exp(-L)]$ (arithmetic mean of passrates) while the loss scaling law gives us $\exp(-\mathbb{E}[L])$ (geometric mean). Using Taylor expansion:
$\mathbb{E}[\exp(-L)]\approx\exp(-\mu_{L})\ +\frac{\sigma_{L}^{2}}{2\mu_{L}}$
where $\mu_{L}$ and $\sigma_{L}^{2}$ are the mean and variance of the loss distribution.
This approximation is accurate when tasks have similar difficulty feature ($\sigma_{L}^{2}/\mu_{L}^{2}\ll 1$), motivating our clustering approach to reduce intra-cluster variance. Assumption 3 adds the parameter $g$ for random guessing. The complete proofs are provided in Appendix B. ∎
theorem˜1 demonstrates that a metric of an evaluation set with similar difficulty features can be effectively modeled using the following formula:
@@
where $a$ and $b$ jointly influence how accuracy varies with $C$, $c$ controls the upper bound of the fitting curve, and $g$ represents the expected random guess metric for a task cluster. $a$, $b$, $c$, and $g$ are trainable parameters. Note that these assumptions may not perfectly hold in practice, we provide additional discussions on assumption 3 in Appendix I.
@@ 4.3 Extrapolation
To ensure reliable extrapolation, we identify clusters exhibiting robust scaling patterns, as some clusters may show saturated or non-emergent performance on smaller models, making them unsuitable for prediction. We aim to find an extrapolation subsets that represent the full set performance, and use the subset metric as a intermediate indicator for the prediction of the full set accuracy.
@@
We filter out non-extrapolatable clusters using two rules based on the parameters from Eq.˜4:
@@
In practice, for extrapolatable clusters, we set $a>1$, $b>0.1$, and $0\leq c<1$. Further ablation experiments are provided in Appendix C.1.
@@ 4.4 Mapping from Predictable Subset to Target Evaluation Set
We map predictions from the predictable subset $\mathcal{P}^{\prime}$ to the complete evaluation set $\mathcal{P}$ using a function $g$. This approach is based on the observation that extrapolatable and non-extrapolatable samples, while differing in difficulty, often share question types, suggesting a preserved relative metric ordering. The mapping function $f:\text{Acc}(\mathcal{P}^{\prime})\rightarrow\text{Acc}(\mathcal{P})$ is continuous, smooth over $[0,1]$, monotonically increasing, and constrained to pass through $(0,0)$ and $(1,1)$. Empirical validation indicates a polynomial function optimally captures this relationship:
$f_{n}(x)=\alpha_{n-1}x^{n}+\alpha_{n-2}x^{n-1}+...+\alpha_{1}x^{2}+(1-\alpha_{1}-\alpha_{2}-...-\alpha_{n-1})x,$ (5)
where $x$ represents the average accuracy of the predictable subset $\mathcal{P}^{\prime}$, and $n$ denotes the highest degree of the polynomial. In practical implementation, a quartic function is employed to fit the mapping (see Appendix C.2 for exhaustive experimental details). To ensure reliability, we calibrate $f$ using evaluation results from existing models as anchors. This subset-to-full mapping generally demonstrates robustness across diverse model architectures and training data, often permitting the use of external models (e.g., Qwen2-72B [45]) as anchors for many tasks (see Appendix C.3 for experiments). However, for data-sensitive tasks, models with similar training distributions provide more reliable anchors, prioritizing data consistency for mapping accuracy. The final metric prediction for a target LLM with estimated training computation $C_{0}$ is then $p=f\circ y(C_{0})$, combining the cluster-wise extrapolation $y(C_{0})$ from Eq.˜4 with the mapping $f$.
@@ 5.1 Experimental Setups
In our experimental setup, we train nine language models ranging from 122M to 70B parameters in total, which share the same data distribution and architecture, with the training data scaled proportionally to their sizes. We show the detailed training configurations and recip in Appendix D.
For evaluation, we adopt the following widely used benchmarks, including GSM8K [9], MATH [19], BBH [38], TriviaQA [24], MBPP [4], AGIEval [50], DROP [12], MMLU-pro [41]. All models are evaluated in a few-shot in-context learning manner, and we aligned our evaluation setups with LLaMa3 [13].We evaluate our proposed COD performance scaling for LLMs against existing approaches on multiple public benchmarks. Using eight smaller language models as known information, we estimate the downstream task performance of a pretrained LLM with 70B parameters.